A multiple choice test has 30 questions. Each question has five possible answers, of which one is correct. If a student guesses on every question, find the probability of getting exactly 12 correct.
Answer by stanbon(48568)

A multiple choice test has 30 questions. Each question has five possible answers, of which one is correct. If a student guesses on every question, find the probability of getting exactly 12 correct.
-------
P(correct answer)=1/5 = 0.2
Your problem is binomial with n=30, p=0.2, X=12
Prob(X=12)=30C12(0.02)^12*(.08)^18=0.006
Cheers,
Stan H.